STScI Logo

Hubble Space Telescope
NICMOS STScI Analysis Newsletter 38

| STScI Analysis Newsletter
| 07 September 2007
Visit the NICMOS web site "New in  
the Last 45 Days" for all new information about NICMOS.


-NICMOS Safing and Recovery
-A new Instrument Science Report regarding NICMOS Time Dependent  
-A new Instrument Science Report regarding NICMOS Focus Update
-A new Instrument Science Report regarding NICMOS Cycle 15 Baseline  
Calibration Plans
-A new Instrument Science Report regarding Corrections to a Subset  
of NICMOS Parallel Aperture Locations


NICMOS Safing and Recovery

NICMOS entered safe mode on Saturday, September 1, 2007 at 10:58 PM  
EDT (Day 245, 02:58 UTC), triggered by an apparent Analog/Digital  
FIFO read failure of internal engineering status by the Control  
Section flight software of the NICMOS main computer. The most  
likely candidate for the cause of this error was a transient single  
event upset affecting either the special engineering data buffer, a  
CPU register, or a memory location.  This event occurred during  
orbit day outside the South Atlantic Anomaly.  HST was in two-gyro  
mode following the successful recovery from the failure of Gyro #2  
on the previous day, which appears unrelated to the NICMOS safing,  
and NICMOS was not carrying out any observations at the time.  All  
remaining NICMOS observations for the week of September 3, 2007  
were removed from the calendar until the instrument could be  

All other NICMOS telemetry diagnostics remained nominal up to the  
point at which the error occured, and there is no indication of any  
failure in the NICMOS electronics or mechanisms. The NCS was also  
not affected and continued operating nominally during the event.  
Similar single event upsets have been observed in the past on other  
instruments as transient events. As no credible risk to the  
instrument could be identified, HST Project proceeded with the  
recovery procedure. NICMOS was successfully transitioned from safe  
to SAA operate on September 4, 2007 at 1:10 PM EDT (Day 247, 17:10  
UTC), and by September 6, 2007 all three internal buffer boxes had  
reached their nominal operating temperatures. No reoccurence of the  
error has been observed. NICMOS science observations are scheduled  
to resume as normal during the week starting September 10, 2007.


NICMOS ISR 2007-002:  NICMOS Time Dependent Flat-fields
T. Dahlen, H. McLaughlin and R. de Jong

We use data obtained during the NICMOS flat-field monitoring  
2002-2007 to derive a set of time dependent flat-fields. These  
epoch flat-fields are compared to the current pipeline flat-fields  
and differences are quantified. Our results are consistent with a  
decrease in DQE over the five-year period on the order of one per  
cent. We also find that the spatial structure of the flat-fields  
change with time. The reasons for these changes are not fully  
known, but we note that the changes are consistent with a change in  
detector temperature with time. The order of this effect is  
generally less than one percent.


NICMOS ISR 2007-003:  NICMOS Focus Update
H. McLaughlin and T. Wiklind

This ISR addresses a number of improvements in the way the NICMOS  
focus is monitored. It provides a detailed explanation of those  
changes and a step by step description of the new procedure. The  
ISR also highlights the results of the focus checks after the  
January 2007 HST safing event, which caused the NICMOS cooling  
system (NCS) anomalous shutdown in February of 2007.


NICMOS ISR 2007-004:  NICMOS Cycle 15 Baseline Calibration Plans
N. Pirzkal, E. Barker, L. Bergeron, R. Bohlin, T. Dahlen, R. De  
Jong, A. M. Koekemoer, H. McLaughlin, B. Shaw, T. Wiklind

This document summarizes the NICMOS Calibration Plans for Cycle  
15.  The monitoring programs from the previous Cycle 13 and 14 are  
carried out in Cycle 15. Additional calibration is included in  
order to improve the accuracy of previously archived NICMOS data.


NICMOS TIR 2007-001:  Corrections to a Subset of NICMOS Parallel  
Aperture Locations
M. Lallo, C. Cox, I. Dashevsky, A. M. Koekemoer

While delivering recently required new SIAF entries representing  
defocused NICMOS apertures, we discovered incorrect values in  
existence for a specific set of NICMOS parallel camera positions.  
We investigated and understood the cause, determined the correct  
values, proposed a fix, both retroactive and for the future then  
delivered the corrected data to operations. The error, which was  
extant for at least five years (since cryo-cooler operations), was  
unnoticed by GOs, the IDT, and the STScI due to the fact that it  
affected only the astrometry of certain NICMOS camera FOVs when  
they were used as parallel with a PAM setting that placed them  
significantly out of focus (~0.5 micron rms wavefront error).


For more detailed information and updates, please consult the
NICMOS webpages:

| Help:
| To subscribe or unsubscribe to the STAN, send a message
| to with the subject line blank and
| the following in the body: [un]subscribe nicmos_news
| The Space Telescope Science Institute is operated by the
| Association of Universities for Research in Astronomy,
| Inc., under NASA contract NAS5-26555.